# LLM Analysis of 150+ years of German Parliamentary Debates on Migration Reveals Shift from Post-War Solidarity to Anti-Solidarity in the Last Decade

### v1

@@ Appendix C Models Inference and Training Details
Figure 17: High-level confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined. Human Annotation scores are computed by averaging the confusion matrices from four leave-one-out (LOO) comparisons, where each annotator is compared to the consensus of the others.
@@
Figure 18: Fine-grained confusion matrices showing the best-performing configuration for each model on the two migrant test sets (Test 1 and Test 2), combined. Human Annotation scores are computed by averaging the confusion matrices from four leave-one-out (LOO) comparisons, where each annotator is compared to the consensus of the others.
@@
Table 3: F1 scores per model for fine-grained and high-level solidarity classification for Migrant (both test sets combined) and Woman. The top section contains fine-grained categories, and the bold rows show performance on high-level classification. Scores are chosen for the best performing configuration for the model in terms of Macro F1.
Label
GPT-4
Mistral
2-Large
Llama
3.3-70B
Llama
3-8B
Gemma
2-27B
Gemma
2-9B
BERT SBERT
Group-based Solidarity 0.33 0.30 0.29 0.25 0.25 0.34 0.32 0.16
Exchange-based Solidarity 0.56 0.39 0.36 0.25 0.41 0.20 0.21 0.19
Empathic Solidarity 0.04 0.27 0.10 0.08 0.15 0.14 0.00 0.03
Compassionate Solidarity 0.48 0.42 0.51 0.22 0.45 0.43 0.51 0.41
Solidarity (high-level) 0.78 0.52 0.69 0.60 0.58 0.63 0.66 0.51
Group-based Anti-solidarity 0.47 0.25 0.42 0.32 0.42 0.41 0.28 0.20
Exchange-based Anti-solidarity 0.46 0.29 0.35 0.12 0.28 0.29 0.09 0.22
Empathic Anti-solidarity 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Compassionate Anti-solidarity 0.00 0.18 0.07 0.00 0.19 0.11 0.17 0.12
Anti-solidarity (high-level) 0.64 0.53 0.52 0.44 0.56 0.54 0.35 0.35
Mixed 0.38 0.26 0.43 0.10 0.17 0.19 0.24 0.15
None 0.67 0.60 0.63 0.56 0.57 0.58 0.59 0.51
Label
GPT-4
Mistral
2-Large
Llama
3.3-70B
Llama
3-8B
Gemma
2-27B
Gemma
2-9B
BERT SBERT
Group-based Solidarity 0.28 0.24 0.25 0.00 0.20 0.20 0.24 0.08
Exchange-based Solidarity 0.31 0.00 0.00 0.00 0.22 0.30 0.35 0.15
Empathic Solidarity 0.27 0.36 0.41 0.24 0.33 0.41 0.22 0.14
Compassionate Solidarity 0.43 0.32 0.51 0.00 0.37 0.21 0.62 0.52
Solidarity (high-level) 0.85 0.67 0.79 0.51 0.75 0.70 0.77 0.78
Group-based Anti-solidarity 0.31 0.36 0.22 0.00 0.27 0.26 0.00 0.00
Exchange-based Anti-solidarity 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Empathic Anti-solidarity 0.67 0.00 0.25 0.07 0.00 0.17 0.00 0.00
Compassionate Anti-solidarity 0.50 0.33 0.22 0.36 0.15 0.00 0.00 0.00
Anti-solidarity (high-level) 0.65 0.44 0.39 0.31 0.48 0.35 0.12 0.10
Mixed 0.30 0.40 0.37 0.16 0.32 0.30 0.15 0.26
None 0.62 0.54 0.47 0.41 0.66 0.41 0.21 0.40
Table 4: Agreement between GPT-4, Llama-3.3-70B, and human annotations on the Migrant dataset. Subtable LABEL:tab:model-human-agreement shows instance-level agreement; Subtable LABEL:tab:trend-agreement reports trend-level correlation across decades.
Comparison Cohen’s $\kappa$ / Macro F1 N
GPT-4 vs Llama-3.3-70B 0.47 (0.58) / — 17,696
GPT-4 vs Human 0.38 (0.57) / 0.34 (0.63) 2,084
Llama-3.3-70B vs Human 0.36 (0.50) / 0.35 (0.61) 2,084
Label Pearson (p)
Group-based Solidarity 0.560 ($p=2.41\times 10^{-2}$)
Exchange-based Solidarity 0.777 ($p=4.00\times 10^{-4}$)
Empathic Solidarity 0.108 ($p=0.691$)
Compassionate Solidarity 0.915 ($p=7.10\times 10^{-7}$)
Solidarity (high-level) 0.966 ($\boldsymbol{p=1.25\times 10^{-9}}$)
Group-based Anti-solidarity 0.945 ($p=3.65\times 10^{-8}$)
Exchange-based Anti-solidarity 0.898 ($p=2.29\times 10^{-6}$)
Empathic Anti-solidarity 0.160-0.160 ($p=0.553$)
Compassionate Anti-solidarity 0.841 ($p=4.44\times 10^{-5}$)
Anti-solidarity (high-level) 0.976 ($\boldsymbol{p=1.15\times 10^{-10}}$)
Mixed 0.685 ($p=0.0034$)
@@
Figure 20: Percentage of sentences showing solidarity/anti-solidarity per decade for all Migrant keywords from 1867 to 2025. The keywords are sorted by frequency, which means that the reliability decreases towards the bottom-right.
@@
Figure 23: (Anti-)solidarity trends in 1867-2022, restricted to the keywords Flüchtlinge and Ausländer. Figure 23(a) shows the fraction of solidarity (left) and anti-solidarity (right) subtypes predicted by Llama-3.3-70B. Percentages are normalized within each decade across all eight subtypes, so the curves in both panels together sum to 100% per decade. Figure 23(b) shows the aggregated proportions of solidarity, anti-solidarity, and mixed labels per decade; totals may not reach 100% because instances classified as ?None? are omitted (though included in the calculation). The period from 1933 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
@@
Figure 24: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B. The model is first given the high-level classification prompt in Figure 27(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Subfigure 28(a) or Subfigure 29(a) is used.
@@
Figure 25: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Fig. 24).
@@
Figure 26: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Fig. 24).
@@
Figure 27: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset. The model is first given the high-level classification prompt in Figure 27(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Subfigure 28(a) or Subfigure 29(a) is used.
@@
Figure 28: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Fig. 27).
@@
Figure 29: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Fig. 27).
Keyword Solidarity Anti-Solidarity Mixed None Total
Asylsuchende 10 0 0 5 15
Ausländer 168 52 9 525 754
Aussiedler 99 1 5 61 166
Einwanderer 23 1 0 21 45
Emigranten 66 5 0 18 89
Flüchtlinge 1828 23 84 1179 3114
Heimatvertriebene 2471 20 68 1814 4373
Sowjetzonenflüchtlinge 264 5 10 178 457
Zuwanderer 32 0 2 46 80
Zuwanderung 7 0 1 29 37
Table 5: Distribution of Solidarity, Anti-Solidarity, Mixed, and None labels per keyword included in the post-war trends analysis (Section 4.2).
BP CDU/CSU DP DRP DZP FDP GB/BHE KPD SPD WAV Z
Group-based S. 95 707 56 13 23 221 118 50 667 93 5
Exchange-based S. 5 163 5 4 1 46 14 8 73 13 0
Compassionate S. 64 670 38 9 8 135 92 69 537 72 7
Empathic S. 6 91 8 1 0 39 29 4 89 8 1
Solidarity with no subtype 1 18 0 0 0 6 2 1 6 2 0
Group-based AS 28 100 9 3 1 22 9 30 68 23 0
Exchange-based AS 4 14 3 0 0 1 3 3 11 2 0
Compassionate AS 1 3 1 0 0 0 0 3 3 0 0
Empathic AS 0 1 0 0 0 0 0 0 0 0 0
Mixed 15 63 5 1 2 23 8 4 33 3 0
None 81 1564 78 10 11 389 94 54 754 47 4
Total per party 300 3394 203 41 46 882 369 226 2241 263 17
Table 6: Distribution of (anti-)solidarity subtypes by party included in the post-war trends analysis (Section 4.2). AS stands for Anti-Solidarity; S for Solidarity.

### v2

@@ Appendix E Label Distribution Distance Metrics
Figure 19: High-level row-normalized confusion matrices (%) for the best-performing configuration of each model on the two migrant test sets (Test 1 and Test 2, combined). Each row sums to 100, showing how items with a given reference label are distributed across predicted labels. Human Annotators (LOO) shows aggregated leave-one-out annotator comparisons, where each annotator is compared to the consensus of the remaining annotators. The Ensemble model aggregates predictions from Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
@@
Figure 20: Fine-grained row-normalized confusion matrices (%) showing the best-performing configuration for each of the selected models on the two migrant test sets (Test 1 and Test 2), combined. Each row sums to 100, showing how items with a given reference label are distributed across predicted labels. Human Annotators (LOO) shows aggregated leave-one-out annotator comparisons, where each annotator is compared to the consensus of the remaining annotators. The Ensemble model aggregates predictions from Llama-3.3-70B, Qwen-2.5-72B, and gpt-oss-120B.
@@
Table 7: F1 scores for selected models on fine-grained and high-level solidarity classification for Migrant (both test sets combined) and Woman. The top section contains fine-grained categories, and the bold rows show performance on high-level classification. Scores are chosen for the best performing configuration for the model in terms of Macro F1.
Label GPT-5 GPT-4 gpt-oss 120b Qwen 2.5-72B Llama 3.3-70B Llama 4-Scout Mistral 2-Large Gemma 2-27B BERT
Group-based S 0.42 0.33 0.37 0.39 0.29 0.35 0.31 0.25 0.31
Exchange-based S 0.43 0.60 0.55 0.42 0.39 0.44 0.38 0.43 0.22
Empathic S 0.48 0.05 0.26 0.26 0.11 0.17 0.30 0.11 0.00
Compassionate S 0.60 0.50 0.48 0.50 0.57 0.59 0.42 0.46 0.53
Solidarity (high-level) 0.81 0.80 0.78 0.69 0.71 0.78 0.54 0.58 0.67
Group-based AS 0.47 0.51 0.46 0.39 0.47 0.42 0.31 0.44 0.32
Exchange-based AS 0.43 0.48 0.47 0.40 0.42 0.48 0.30 0.30 0.10
Empathic AS 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Compassionate AS 0.20 0.00 0.06 0.16 0.08 0.19 0.16 0.21 0.18
Anti-solidarity (high-level) 0.70 0.69 0.66 0.59 0.56 0.63 0.57 0.58 0.39
Mixed 0.32 0.46 0.32 0.24 0.48 0.08 0.26 0.17 0.26
None 0.66 0.71 0.68 0.65 0.66 0.68 0.64 0.59 0.62
Label GPT-5 GPT-4 gpt-oss 120b Qwen 2.5-72B Llama 3.3-70B Llama 4-Scout Mistral 2-Large Gemma 2-27B BERT
Group-based S 0.45 0.28 0.40 0.41 0.25 0.25 0.24 0.20 0.24
Exchange-based S 0.50 0.31 0.40 0.00 0.00 0.00 0.00 0.22 0.35
Empathic S 0.27 0.27 0.34 0.46 0.41 0.31 0.36 0.33 0.22
Compassionate S 0.62 0.43 0.59 0.42 0.51 0.34 0.32 0.37 0.62
Solidarity (high-level) 0.86 0.85 0.82 0.76 0.79 0.76 0.67 0.75 0.77
Group-based AS 0.22 0.31 0.25 0.27 0.22 0.22 0.36 0.27 0.00
Exchange-based AS 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Empathic AS 0.50 0.67 0.47 0.00 0.25 0.36 0.00 0.00 0.00
Compassionate AS 0.40 0.50 0.50 0.24 0.22 0.30 0.33 0.15 0.00
Anti-solidarity (high-level) 0.59 0.65 0.53 0.30 0.38 0.39 0.44 0.48 0.12
Mixed 0.50 0.30 0.38 0.33 0.37 0.43 0.40 0.32 0.15
None 0.65 0.62 0.48 0.51 0.47 0.26 0.54 0.66 0.21
@@
Figure 22: (Anti-)solidarity trends for the Migrant category by keyword and decade, 1867-2025. Each panel shows DSL-adjusted decade-level shares of solidarity and anti-solidarity for one keyword (see definition (c) in Section˜3.2.2). The mixed and none categories are included in the four-class distribution but omitted from the visualization. Keywords are sorted by frequency, so estimates become less reliable toward the bottom right.
(a) Human majority vote
(b) Human soft labels
(c) gpt-oss-120B
(d) Model majority vote
(e) Llama-3.3-70B
(f) Qwen-2.5-72B
(g) Model soft labels
(h) DSL (soft labels)
Figure 23: High-level solidarity trends for the Migrant category under selected labeling and correction strategies discussed in Section˜3.2.2 based on the 2k evaluation subset. Trends are computed as label shares within each decade (see definitions in Section˜3.2.2). The period from 1930 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
(a) Human majority vote
(b) Human soft labels
(c) gpt-oss-120B
(d) Model majority vote
(e) Llama-3.3-70B
(f) Qwen-2.5-72B
(g) Model soft labels
(h) DSL (soft labels)
Figure 24: Fine-grained solidarity trends for the Migrant category under selected labeling and correction strategies discussed in Section˜3.2.2 based on the 2k evaluation subset. Trends are computed as label shares within each decade (see definitions in Section˜3.2.2). The period from 1930 to 1949 (grey shaded area) is excluded due to limited data availability during and immediately after the NS dictatorship.
Label Model Pearson Spearman RMSE
High-level trends (Human majority vote)
Solidarity Llama-3.3-70B 0.907 0.894 0.101
DSL (soft labels) 0.910 0.938 0.075
Anti-solidarity Llama-3.3-70B 0.886 0.847 0.065
DSL (soft labels) 0.928 0.921 0.041
Mixed Llama-3.3-70B 0.810 0.830 0.029
DSL (soft labels) 0.860 0.873 0.027
None Llama-3.3-70B 0.910 0.906 0.134
DSL (soft labels) 0.868 0.899 0.074
Average Llama-3.3-70B 0.878 0.869 0.082
DSL (soft labels) 0.892 0.908 0.054
Fine-grained trends (Human majority vote)
Group-based S Llama-3.3-70B 0.412 0.358 0.202
DSL (soft labels) 0.679 0.678 0.062
Exchange-based S Llama-3.3-70B 0.588 0.458 0.073
DSL (soft labels) 0.461 0.508 0.055
Empathic S Llama-3.3-70B 0.763 0.827 0.044
DSL (soft labels) 0.559 0.519 0.022
Compassionate S Llama-3.3-70B 0.823 0.772 0.204
DSL (soft labels) 0.872 0.912 0.080
Group-based AS Llama-3.3-70B 0.855 0.824 0.087
DSL (soft labels) 0.930 0.908 0.058
Exchange-based AS Llama-3.3-70B 0.706 0.642 0.024
DSL (soft labels) 0.728 0.729 0.021
Empathic AS Llama-3.3-70B 0.497 0.435 0.017
DSL (soft labels) 0.365 0.442 0.018
Compassionate AS Llama-3.3-70B 0.447 0.492 0.064
DSL (soft labels) 0.842 0.660 0.036
Average Llama-3.3-70B 0.636 0.601 0.089
DSL (soft labels) 0.680 0.670 0.044
Table 8: Per-label correlation between model predictions and human-annotated decade-level trends. Higher is better for Pearson and Spearman; lower is better for RMSE.
@@
Figure 27: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B. The model is first given the high-level classification prompt in Figure˜30(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Figure˜31(a) or Figure˜32(a) is used.
@@
Figure 28: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Figure˜27).
@@
Figure 29: Two-step few-shot prompt set used for the evaluation of open-weight models on the Migrant dataset and for the final large-scale annotation with Llama-3.3-70B (continued from Figure˜27).
@@
Figure 30: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset. The model is first given the high-level classification prompt in Figure˜30(a). If the output is Solidarity or Anti-solidarity, the corresponding subtype prompt in Figure˜31(a) or Figure˜32(a) is used.
@@
Figure 31: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Figure˜30).
@@
Figure 32: Two-step few-shot prompt set used for the evaluation of open-weight models on the Frau dataset (continued from Figure˜30).
Figure 33: Solidarity trends for Migrant speeches computed using DSL-corrected soft labels (see definition (c) in Section˜3.2.2) on the restricted corpus containing only the keyword groups Flüchtlinge and Ausländer. Figure˜33(b) shows the high-level (anti-)solidarity trends, while Figure˜33(a) shows the distribution of fine-grained (anti-)solidarity subtypes over time. In the fine-grained panel, percentages are normalized within each decade across the eight solidarity and anti-solidarity subtypes. The period 1930-1949 (grey area) is excluded due to limited data during and after the Nazi dictatorship.
Keyword Solidarity Anti-Solidarity Mixed None N (statements)
Asylsuchende 4.11 2.06 0.48 3.35 10
Ausländer 291.43 118.86 21.43 326.28 758
Aussiedler 118.43 12.48 15.22 25.87 172
Einwanderer 2.84 3.21 0.20 19.75 26
Emigranten 21.14 11.08 2.61 10.18 45
Flüchtlinge 1917.08 321.69 174.29 859.94 3273
Heimatvertriebene 3158.67 422.75 167.72 1006.85 4756
Sowjetzonenflüchtlinge 308.99 27.13 2.63 138.25 477
Zuwanderer 58.24 6.54 0.86 60.36 126
Table 9: Distribution of DSL-adjusted expected probability mass for solidarity, anti-solidarity, mixed, and none per keyword included in the post-war trends analysis Section˜4.2. Values reflect summed DSL-adjusted expected counts. Totals correspond to the number of keyword-coded statements in the analyzed subset.
Party Solidarity Anti-Solidarity Mixed None N (statements)
BP 134.13 54.77 67.51 46.60 303
CDU/CSU 2040.48 229.98 94.93 1050.62 3416
DP 138.21 40.81 -0.54 26.52 205
DRP 34.53 1.47 1.74 3.26 41
FDP 539.79 38.02 56.78 251.41 886
GB/BHE 318.75 27.61 9.03 17.62 373
KPD 151.75 57.43 -1.14 18.96 227
SPD 1595.06 153.56 37.13 478.25 2264
WAV 191.81 19.72 37.12 15.35 264
Zentrum 53.61 3.02 1.35 5.01 63
Table 10: Distribution of DSL-adjusted expected probability mass for solidarity, anti-solidarity, mixed, and none per party included in the post-war trends analysis Section˜4.2. Values reflect summed DSL-corrected expected counts. Totals correspond to the number of party-coded statements in the analyzed subset.